Pencil lead is actually carbon Part A What is the resistance of the 0.750 mm -diameter, 5.10 cm -long lead from a mechanical pencil? Submit Request Answen Provide Feedback Next >
Answer #1

Here ,

resistivity , p = 3.5 *10^-5 Ohm.m

L = 5.1 cm = 0.051 m

Resitance = p * L/area of cross section

Resistance = 3.5 *10^-5 * 0.051/(pi *(0.750 *10^-3/2)^2)

Resistance = 4.04 Ohm

the resistance of the lead is 4.04 Ohm